# On Speeding Up Language Model Evaluation

### v1

@@ 3.2 Algorithm 2 – UCB-E with Low-Rank Factorization (UCB-E-LRF)
Algorithm 2 UCB-E-LRF ($\mathcal{A}_{uel}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,K,T_{0},b,\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$, the LRF oracle $\mathcal{M}$, the rank $r$, # solutions in ensemble $K$, the warm-up budget $T_{0}$, the batch size $b$, the uncertainty scaling $\eta$.
@@
1:  Uniformly sample $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ and the observation matrix $O\in\{0,1\}^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j}+\eta R_{i,j})$
3:  for $t=T_{0},\cdots,T_{0}+[(T-T_{0})/b]$ do
4:     Select: Draw $i_{t}\in\arg\max_{i}B_{i}$; Draw $J_{t}:=\mathrm{Top-}b\left({R_{i_{t},j}}|j\in[m]\right)$.
5:     for $j_{t}\in J_{t}$ do
6:        Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$
@@
8:     Update: $\forall j_{t}\in J_{t}$, $O_{i_{t},j_{t}}\leftarrow 1$; $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,K)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j}+\eta R_{i,j})$.
9:  end for
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})$
Main idea/Inspiration/Motivation.
We find that the scoring matrix $E$ for real-world problem instances can often be approximated well by a low rank matrix where rank $r\ll n,m$. This observation means that it is possible to estimate the score of all method-example pairs from the scores of just a few. Consider if the matrix $E$ were exact rank-$1$, then only $n+m-1$ scores would be needed to exactly recover the full scoring matrix $E$, which is extremely efficient compared to exhaustively evaluating all $n\cdot m$ combinations. Suppose $E^{\rm obs}$ is a partially observed scoring matrix, where the non-zero elements are evaluated pairs with the ground truth score for that pair, and $\hat{E}$ is the estimated scoring matrix, where the non-zero elements are unevaluated pairs with the estimated score. Suppose $O\in\{0,1\}^{n\times m}$ is the observation matrix where $O_{i,j}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been observed. When the scores of unevaluated method-example pairs are estimated accurately, the score estimation $\hat{\mu}_{i}=\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})$, which combines both evaluated pairs and the estimated pairs, has the potential to be more accurate than the estimation by only the evaluated pairs.
The algorithm.
Leveraging the intuition of estimating the unevaluated method-example pairs, we adopt a low rank factorization (LRF) approach and propose our second algorithm UCB-E-LRF ($\mathcal{A}_{\rm uel}$; Algorithm 2). This algorithm builds on UCB-E which dynamically allocates the budget to the promising methods. We estimate the upper confidence bound for each method $f_{i}$ and evaluate an example $x_{j_{t}}$ yet to be evaluated for the method $f_{i_{t}}$ with largest upper confidence bound. However, the specification of this idea in UCB-E-LRF is different, because of our additional intermediate goal – a good estimation $\hat{E}$ for the unevaluated pairs. Suppose we have a scoring matrix estimation oracle $\mathcal{M}$: given $t$ evaluated method-example pairs, the oracle outputs the estimated scoring matrix $\hat{E}$ and the uncertainty matrix $R$ that quantifies the uncertainty for each estimation in $\hat{E}$. Suppose we keep updating observation matrix $O\in\{0,1\}^{n\times m}$ where $O_{i,j}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been evaluated. The different specifications from UCB-E are:
  1. 1.
    Once the method $f_{i_{t}}$ is selected, unlike in UCB-E where the example $x_{j_{t}}$ is uniformly drawn from the unevaluated examples, we pick the example $x_{j_{t}}$ with the largest uncertainty $R_{i_{t},j_{t}}$ from the unevaluated examples, because evaluating this example $x_{j_{t}}$ will result in the largest reduction in uncertainty for the method $f_{i_{t}}$. This allows us to better estimate the scoring matrix $\hat{E}$ with less evaluations.
  2. 2.
    The upper confidence bound is now computed as $\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j}+\eta R_{i,j})$, which is a combination of the current score estimation $\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})$ and the uncertainty estimation $\frac{1}{m}\sum_{j=1}^{m}\eta R_{i,j}$. Notably, the current score estimation is the partially observed scores $\frac{1}{m}\sum_{j=1}^{m}O_{i,j}E^{\rm obs}_{i,j}$ plus the estimated scores $\frac{1}{m}\sum_{j=1}^{m}(1-O_{i,j})\hat{E}_{i,j}$
The choice of scoring matrix estimation oracle.
We now introduce the scoring matrix estimation oracle $\mathcal{M}$ used in this paper. The inputs of the oracle $\mathcal{M}$ are a partially observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ and an observation matrix $O\in\{0,1\}^{n\times m}$ where $O_{i,j}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been evaluated. The outputs of $\mathcal{M}$ are supposed to be an estimated scoring matrix $\hat{E}$ and the corresponding uncertainty matrix $R$.
We define our $\mathcal{M}$ as an ensemble of $K$ low-rank factorization (LRF) solutions [59, 14]. The low-rank factorization assumes the ground truth matrix is (approximately) low-rank with rank $r$. We optimize the low-rank representations $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{m\times r}$ for following optimization problem by the alternating least squares method [26]:
$\min_{U\in\mathbb{R}^{n\times r},V\in\mathbb{R}^{m\times r}}\sum_{i\in[n],j\in[m]}O_{i,j}\left(U_{i}^{\top}V_{j}-E^{\rm obs}_{i,j}\right)^{2},$ (1)
where $U_{i}$ is the $i$th row of $U$ and $V_{j}$ is the $j$th row of $V$. We further produce both the estimation and the uncertainty by bootstrapping [23, 67]. In order to get a diverse ensemble of low-rank factorizations for each factorization $(U^{(k)},V^{(k)})$, we randomly zero out $5\%$ of the “1" in $O$ as $O_{k}$ and optimize the objective Equation 1 with $E^{\rm obs}$ and $O_{k}$ instead. Finally we define the estimated scoring matrix $\hat{E}$ and the uncertainty matrix $R$ as
$\hat{E}:=(\mathbf{1}-O)\circ\left(\frac{1}{K}\sum_{k=1}^{K}U^{(k)}\left(V^{(k)}\right)^{\top}\right),R_{i,j}:=\sqrt{\frac{1}{K}\sum_{k=1}^{K}((U^{(k)}\left(V^{(k)}\right)^{\top})_{i,j}-\hat{E}_{i,j})^{2}}$ (2)
where $\mathbf{1}$ is an all-one matrix with the shape of $n\times m$ and $\circ$ is the element-wise multiplication.
This oracle introduces four additional hyperparameters for our UCB-E-LRF algorithm, which determine performance and efficiency:
  • The rank $r$ of the low-rank factorization. The value of $r$ adjusts the expressiveness of the factorization when fit to the data in Equation 2. Setting $r$ requires considering the bias-variance trade-off of choosing smaller vs. larger $r$. The representations of smaller $r$ inherently introduce larger approximation error, however, the representations of large $r$ are harder to be estimated accurately given only a few observations in $E^{\rm obs}$.
  • The ensemble size $K$. Sufficiently large $K$ is necessary to avoid additional noise when estimating $\hat{E}$ and $R$ in Equation 2. Yet, the computational cost of the factorization scales with $K$ and therefore $K$ cannot be too large.
  • The warm-up budget $T_{0}$. The oracle $\mathcal{M}$ requires a minimum number of initial observations in order to estimate the low-rank representation accurately. Therefore, before the active selection phase (line 3-9 in Algorithm 2), we have a warm-up phase (line 1-2 in Algorithm 2), where $T_{0}$ method-example pairs are sample uniformly at random to evaluate before moving on to the “active" phase.
  • The batch size $b$. The oracle $\mathcal{M}$ must be fit and refit which involves optimizing the ensemble of low-rank factorizations. As the computational cost of this procedure cannot be neglected, we introduce a batch size hyperparameter. At each time step $t$, we select $b$ examples $J_{t}$ paired with the method $i_{t}$ (line 4 in Algorithm 2), evaluate the $b$ method-example pairs (line 6 in Algorithm 2), and update the intermediate variables with these $b$ scores (line 8 in Algorithm 2). Notice that although increasing $b$ is more computationally friendly as it calls the oracle $\mathcal{M}$ less, it may hurt the performance as it reduces the granularity of the decisions, making the algorithm less “active". The choice of $b$ is ultimately a trade-off between computational cost and the performance, however, we will see in the experiments section that performance does not vary drastically for many choices of $b$.
We analyze these four UCB-E-LRF hyperparameters in our experiment sections, along with the uncertainty scaling $\eta$ in both UCB-E and UCB-E-LRF.
@@ 4 Experiments
Figure 4: Ablations of effect of hyperparameters on UCB-E-LRF algorithm. The results suggest that our selected hyperparameters are efficient and effective.
@@ 4.1 Datasets
To assess the performance of our algorithms under a variety of use cases, we test with three datasets AlpacaEval [44], Grade School Math 8K (GSM8K) [18] and Physical Interaction: Question Answering (PIQA), together with different settings of the method set $\mathcal{F}$ and the scoring function $e$; Table 2 summarizes the statistics of each dataset along with what method $\mathcal{F}$ and scoring function $e$. For AlpacaEval, we design two sets of $calF$. The first set $\mathcal{F}$ contains all LLMs reported by  [44]. The second $\mathcal{F}$ contains the same LLMs except for the annotator model GPT-4 Turbo. The latter $\mathcal{F}$ that does not contain GPT-4 Turbo makes the learning more challenging and interesting since the annotator, GPT-4 Turbo, is a clear winner among all LLMs. For GSM8K and PIQA, we set $\mathcal{F}$ as a set of prompts to simulate prompt engineering, or a set of LLMs with various sampling configurations, to mimic model selection and hyperparameter tuning. We provide more information about these datasets in Appendix A.1.
Table 2: Statistics of each dataset used in our experiments. The examples $\mathcal{X}$ of each dataset are the questions from their respective evaluation benchmark. More details of the datasets can be found in Appendix A.1.
@@ 4.2 Metrics
NDCG.
Although our focus is to identify the best method quickly, it is sometimes also desirable to generally rank top-$P$ promising methods high. We therefore evaluate with normalized discounted cumulative gain (NDCG) at $P$. NDCG@$P$ takes as input the top-$P$ prediction by an algorithm and the higher their true ranks are, the higher the NDCG is. We choose $P=10$ to evaluate all algorithms.
@@ 4.3 Set-Up of Our algorithms and Baselines
We introduce the set-up of our two algorithms, UCB-E and UCB-LRF, together with three more baselines, Row Mean Imputation, Filled Subset and LRF.
@@
Our proposed algorithm as shown in Algorithm 1. We use $\eta=1$ since this consistently yields the best performance across all datasets.
@@
Our proposed algorithm as shown in Algorithm 2. For all datasets, we use rank $r=1$ for low rank factorization with an ensemble size of $K=64$. We use $5\%$ of data for warm up i.e. $T_{0}=0.05\times n\times m$ and $\eta=5$. We use a batch size of $b=32$.
@@ 4.4 Main Results
In Figure 3, we plot the performance of baselines and our algorithms on all six datasets (columns) as the budget $T$ increases from 5% to 100% of the total number of method-example pairs. In each row, we evaluate the algorithms on a different metric (either top 1 precision with different $\epsilon$, $p$ or NDCG@10). Each line in the figure is the average result over 50 independent trials with different seeds. For example, an average top 1 precision of 0.9 indicates that 45 out of 50 trials predict the best method correctly at the budget level that its x-coordinate represents. In addition, we calculate $H_{1}$ as defined in Corollary 1 that quantifies the difficulty of finding the best method on a dataset. Intuitively, a higher $H_{1}$ value suggests that the method set $\mathcal{F}$ is large or there are many methods that have similar performance with the best one and distinguishing them can be challenging. In Appendix A.1 Figure 5, we also plot the histogram to show the distribution of performance among $\mathcal{F}$ on these datasets.
@@ 4.5 More Empirical Analysis
To study the effect of selecting both next method and next example to evaluate using uncertainty matrix $R$, we consider an ablation of UCB-E-LRF by reverting the place of using $R$ to the UCB-E: UCB-E-LRF (Score Only) where the upper confidence bound is computed by the confidence interval in original UCB-E in addition to the mean estimation through low rank factorization; the example is uniformly selected, same as UCB-E. The detailed algorithmic description can be found in Algorithm 6. We see that on the hard dataset GSM8K Prompts where UCB-E-LRF has certain benefit over UCB-E, there is a significant gap between UCB-E-LRF and UCB-E-LRF (Score Only). This means that the component of uncertainty matrix $R$ in UCB-E-LRF is crucial to contribute the benefit over UBC-E.
@@
Intuitively, a smaller batch size is more flexible can have more fine grained selection. We experiment with $b\in\{2,8,32,128\}$ and as shown in the plot, smallest batch size gives the best performance especially on harder datasets whereas a batch size of 128 significantly degrades the performance. However, it incurs more overall computational cost due to fitting low rank factorization more often than a smaller batch size; for example when $b=2$, it takes almost 16 times as much time as $b=32$ to achieve $1.0$ precision, which is ineffective. To achieve a balance of performance and computational cost, we use a batch size of $32$ in our experiment.
Ensemble Size $K$ Ablation.
We vary the ensemble size $K\in\{4,16,64,256\}$ and find that very small ensemble size gives suboptimal performance on hard datasets. There is almost no performance difference on easy dataset like AlpacaEval (Drop Annotator). The performance is robust to different $K$ as long as it is larger than $64$.
@@
We experiment with different uncertainty scaling values $\{0.1,0.5,1,5,10\}$. In Figure 4, it can be seen that a certain range of $\eta$, from $0.1$ to $5$, gives similar performance on both two datasets. This demonstrate the robustness of selecting $\eta$.
@@
As discussed in the Algorithm Section, $r$ adjusts the bias-variance trade-off. Empirically we experiment with $r\in\{1,2,5\}$ and find that $r=1$ is consistently better than larger value counterparts. The results can be explained by the fact that a larger $r$ requires more evaluated data in order to prevent overfitting, which might not have advantage in the limited budget setting. Therefore, we recommend using $r=1$ for all datasets.

### v2

@@ 3.2 Algorithm 2 – UCB-E with Low-Rank Factorization (UCB-E-LRF)
Algorithm 1, the UCB-E algorithm, treats each method independently and samples examples for a method uniformly at random. In practice, however, methods and examples are correlated with each other and we can predict the outcome of a particular method-example pair from previous observations. Being able to accurately predict the outcome gives better estimate of the score of a method. Additionally, instead of sampling examples uniformly at random, prioritizing examples with large prediction uncertainties reduces the overall uncertainty in the scores of each method, which provides more accurate score estimates in the subsequent steps. In this section, we propose a low-rank factorization based algorithm (UCB-E-LRF) to capture the above intuition.
Low-rank factorization.
A natural approach to modeling these method-example interactions is through low-rank matrix factorization. Intuitively, if the method-examples are very correlated, there should exist $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{m\times r}$ and rank $r\ll\min(n,m)$ such that $E\approx UV^{\top}$. Obtaining $U$ and $V$ from observed scoring matrix $E^{\rm obs}$ allows us to estimate the scores of all method-example pairs from just a few scores. For instance, if the matrix $E$ were an exact rank-1 matrix, only $n+m-1$ scores would be needed to recover the full scoring matrix $E$, which is far more efficient than exhaustively evaluating all $n\times m$ combinations. To find $U$ and $V$, we use alternating least squares [26] to optimize the following problem:
$\min_{U,V}\Big{|}\Big{|}O\odot\Big{(}UV^{T}-E^{\rm obs}\Big{)}\Big{|}\Big{|}_{F}^{2}.$ (1)
Score estimation.
We can then combine the observed matrix $E^{\rm obs}$ and estimated matrix $\hat{E}=UV^{T}$, using $O$ as a gating function, to arrive at a prediction of the score $\mu_{i}$ as follows:
$\hat{\mu}_{i}=\frac{1}{m}\sum_{j=1}^{m}O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j}.$ (2)
Uncertainty estimation.
The score estimation gives us a better way to estimate score of every method by exploiting the correlation between methods and examples at every step. We now discuss how we estimate prediction uncertainties with ensembles.
By ensembling the factorization model, we obtain an empirical distribution of predictions for each method-example pair. Consequently, the standard deviation of the predictions can be used to quantify uncertainty. If the predictions are very similar to each other, this suggests the ensembles tend to agree with each other, and hence there is little uncertainty. In order to achieve sufficient model variance to approximate uncertainty, we uniformly at random drop some values from $E^{\rm obs}$ when solving the factorization problem in Equation 1 for each model  [23, 67]. For an ensemble of size $S$ the estimate for $\hat{E}$ becomes the following:
$\hat{E}=\frac{1}{S}\sum_{s=1}^{S}U^{(s)}(V^{(s)})^{T}.$ (3)
Then the uncertainty matrix $R\in\mathbb{R}^{n\times m}$ is given by:
$R=\sqrt{\frac{1}{S}\sum_{s=1}^{S}\Bigg{[}\Big{(}1-O\Big{)}\odot\Big{(}U^{(s)}(V^{(s)})^{T}-\hat{E}\Big{)}\Bigg{]}^{2}}.$ (4)
Adapative selection.
Bringing score estimation and uncertainty estimation together, we propose our second algorithm UCB-E-LRF ($\mathcal{A}_{\rm uel}$; Algorithm 2), an extension of Algorithm 1 leveraging the low-rank factorization model $\mathcal{M}$. The high level structure of the algorithms are very similar to each other. Specifically, we estimate the upper confidence bound (UCB) for each method based on a linear combination of score and uncertainty estimation, select the method with the greatest UCB, and evaluate the method on an example that has the highest uncertainty. Lastly, we introduce a warm-up budget $T_{0}$. This is because $\mathcal{M}$ requires a minimum number of initial observations in order to estimate the low-rank representation accurately. Therefore, before the active selection phase (line 3-9 in Algorithm 2), we have a warm-up phase (line 1-2 in Algorithm 2), where $T_{0}$ method-example pairs are sample uniformly at random to evaluate before moving on to the “active" phase.
Algorithm 2 UCB-E-LRF ($\mathcal{A}_{\rm uel}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$
3:  for $t=T_{0},\cdots,T$ do
4:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\arg\max_{k\,\mid\,O_{ik}=0}R_{ik}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
6:     Update: $O_{i,j}\leftarrow 1$; $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij})$
@@ 4 Experiments
In this section, we study the performance of UCB-E and UCB-E-LRF in a variety of real settings. We begin by describing our dataset and metrics, and then analyze the experimental results and ablations.
@@
Figure 4: Ablations of the proposed algorithms and hyperparameters. UCB-E and UCB-E-LRF are denoted by blue and red lines, respectively, consistent with Figure 3.
@@ 4.1 Datasets
To assess the performance of our algorithms under a variety of use cases, we test with three datasets AlpacaEval [44], Grade School Math 8K (GSM8K) [18] and Physical Interaction: Question Answering (PIQA) [3], together with different settings of the method set $\mathcal{F}$ and the scoring function $e$; Table 2 summarizes the statistics of each dataset along with what method $\mathcal{F}$ and scoring function $e$. For AlpacaEval, we design two sets of $\mathcal{F}$. The first set $\mathcal{F}$ contains all LLMs reported by [44]. The second $\mathcal{F}$ contains the same LLMs except for the annotator model GPT-4 Turbo. The latter $\mathcal{F}$ that does not contain GPT-4 Turbo makes the learning more challenging and interesting since the annotator, GPT-4 Turbo, is a clear winner among all LLMs. For GSM8K and PIQA, we set $\mathcal{F}$ as a set of prompts to simulate prompt engineering, or a set of LLMs with various sampling configurations, to mimic model selection and hyperparameter tuning. We provide more information about these datasets in Appendix A.
Table 2: Statistics of each dataset used in our experiments. The examples $\mathcal{X}$ of each dataset are the questions from their respective evaluation benchmark. More details of the datasets can be found in Appendix A.
@@ 4.2 Metrics
NDCG@K.
Although our focus is to identify the best method quickly, it is sometimes also desirable to generally rank top K promising methods high. We therefore evaluate with normalized discounted cumulative gain (NDCG) at K. NDCG@K takes as input the top K prediction by an algorithm and the higher their true ranks are, the higher the NDCG is. We choose $K=10$ to evaluate all algorithms.
@@ 4.3 Set-Up of Our algorithms and Baselines
We introduce the set-up of our two algorithms, UCB-E and UCB-LRF, together with three baselines, Row Mean Imputation, Filled Subset and LRF.
@@
Our proposed algorithm as shown in Algorithm 1. We use $a=1$ since this consistently yields the best performance across all datasets. This parameter is ablated in Figure 7 of Appendix B.
@@
Our proposed algorithm as shown in Algorithm 2. For all datasets, we use rank $r=1$ for low rank factorization with an ensemble size of $S=64$. We use $5\%$ of data for warm up i.e. $T_{0}=0.05\times n\times m$ and $\eta=5$.
@@
Finally, for better practicality and efficiency, all algorithms and baselines are implemented with a batch size $b$. That is, instead of selecting and evaluating just one example at a time, the algorithms select $b$ examples. For algorithms that involve randomly drawing an example, they are modified to randomly drawing $b$ examples without replacement. For algorithms that draw an example with the $\arg\max$ operator, they are modified to drawing with the $\mathrm{Top-}b$ operator. We use $b=32$ for reporting all main results.
@@ 4.4 Main Results
In Figure 3, we plot the performance of baselines and our algorithms on all six datasets (columns) as the budget $T$ increases from 5% to 100% of the total number of method-example pairs. In each row, we evaluate the algorithms on a different metric (either top 1 precision with different $\epsilon$, $p$ or NDCG@10). Each line in the figure is the average result over 50 independent trials with different seeds. For example, an average top 1 precision of 0.9 indicates that 45 out of 50 trials predict the best method correctly at the budget level that its x-coordinate represents. In addition, we calculate $H_{1}$ as defined in Corollary 1 that quantifies the difficulty of finding the best method on a dataset. Intuitively, a higher $H_{1}$ value suggests that the method set $\mathcal{F}$ is large or there are many methods that have similar performance with the best one and distinguishing them can be challenging. In Appendix A Figure 5, we also plot the histogram to show the distribution of performance among $\mathcal{F}$ on these datasets.
@@ 4.5 More Empirical Analysis
To study the effect of selecting both the next method and the next example to evaluate using the uncertainty matrix $R$, we consider an ablation of UCB-E-LRF. In this variant, called UCB-E-LRF (Score Only), we modify the UCB-E algorithm by replacing its mean calculation with the mean calculated via low-rank factorization (as described in Equation 3). The upper confidence bound calculation and next example selection is the same as the original UCB-E algorithm.
The detailed algorithmic description can be found in Algorithm 6. We see that on the hard dataset GSM8K Prompts where UCB-E-LRF has certain benefit over UCB-E, there is a significant gap between UCB-E-LRF and UCB-E-LRF (Score Only). This means that the component of uncertainty matrix $R$ in UCB-E-LRF is crucial to contribute the benefit over UBC-E.
@@
Intuitively, a smaller batch size is more flexible can have more fine grained selection. We experiment with $b\in\{1,32,128\}$ and as shown in the plot, smallest batch size gives the best performance especially on harder datasets whereas a batch size of 128 significantly degrades the performance. However, it incurs more overall computational cost due to fitting low rank factorization more often than a smaller batch size; for example when $b=2$, it takes almost 16 times as much time as $b=32$ to achieve $1.0$ precision, which is ineffective.
Ensemble Size $S$ Ablation.
We vary the ensemble size $S\in\{4,16,64,256\}$ and find that very small ensemble size gives suboptimal performance on hard datasets. There is almost no performance difference on easy dataset like AlpacaEval (Drop Annotator). The performance is robust to different $S$ as long as it is larger than $64$.
@@
We experiment with different uncertainty scaling values $\{0.1,0.5,1,5,10\}$. In Figure 4d, it can be seen that a certain range of $\eta$, from $0.1$ to $5$, gives similar performance on both two datasets. This demonstrate the robustness of selecting $\eta$.
@@
Hyperparameter $r$ adjusts the bias-variance trade-off. Empirically we experiment with $r\in\{1,2,5\}$ and find that $r=1$ is consistently better than larger value counterparts. The results can be explained by the fact that a larger $r$ requires more evaluated data in order to prevent overfitting, which might not have advantage in the limited budget setting. Therefore, we recommend using $r=1$ for all datasets.